### v1

# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

Abstract
Superalignment, where humans are weak supervisors of superhuman models, has become an important and widely discussed issue in the current era of rapid development of Large Language Models (LLMs). The recent work preliminarily studies this problem by using weak models to supervise strong models. It discovers that weakly supervised strong students can consistently outperform weak teachers towards the alignment target, leading to a weak-to-strong generalization phenomenon. However, we are concerned that behind such a promising phenomenon, whether there exists an issue of weak-to-strong deception, where strong models may deceive weak models by exhibiting well-aligned in areas known to weak models but producing misaligned behaviors in cases weak models do not know. We then take an initial step towards exploring this security issue in a specific but realistic multi-objective alignment case, where there may be some alignment targets conflicting with each other (e.g., helpfulness v.s. harmlessness). Such a conflict is likely to cause strong models to deceive weak models in one alignment dimension to gain high reward in other alignment dimension. Our experiments on both the reward modeling task and the preference optimization scenario indicate: (1) the weak-to-strong deception exists; (2) the deception phenomenon may intensify as the capability gap between weak and strong models increases. We also discuss potential solutions and find bootstrapping with an intermediate model can mitigate the deception to some extent. Our work highlights the urgent need to pay more attention to the true reliability of superalignment. [^1] Code is available at https://github.com/keven980716/weak-to-strong-deception.
Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization
Wenkai Yang1, Shiqi Shen2, Guangyao Shen2, Zhi Gong2, Yankai Lin1thanks: Corresponding Author 1Gaoling School of Artificial Intelligence, Renmin University of China, Beijing, China 2WeChat, Tencent Inc., Beijing, China {wenkaiyang, yankailin}@ruc.edu.cn
1 Introduction
Human supervision is an indispensable part of the process of constructing practical Large Language Models (LLMs) (Touvron et al., 2023; MetaAI, 2024). Human-annotated data is not only commonly used to enable LLMs to learn human knowledge and accomplish real-world tasks (Wang et al., 2018; Wei et al., 2021; Longpre et al., 2023), but also crucial for aligning models’ behavior with human values Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022); Rafailov et al. (2024).
The recent significant advancements of LLMs (OpenAI, 2022, 2023) suggest that in the near future, LLMs may become superhuman models that are more knowledgeable and intelligent than humans (Burns et al., 2023). In such a superalignment case where humans now become weak supervisors (refer to Figure 1 (a)), it is crucial to study whether supermodels trained under weak human data can demonstrate full potential and most importantly, still align well with human values. Though studying the above problem is intractable today, Burns et al. (2023) take a preliminary step to study in an analogous setting (refer to Figure 1 (b)), where weak language models (e.g., GPT-2 (Radford et al., 2019)) are used to supervise strong language models (e.g., GPT-4 (OpenAI, 2023)). It has been found that the weak supervision can effectively unleash the capabilities of strong models and enable strong models to exhibit better performance than weak teachers. It is called the weak-to-strong generalization phenomenon (refer to Figure 1 (c)).
Figure 1: Illustrations of the concepts discussed in this paper. Importantly, we aim to explore behind the current promising weak-to-strong generalization phenomenon, whether the strong student will selectively exhibit misalignment in the areas of knowledge that are unknown to the weak supervisor. We preliminarily study this problem in a realistic multi-objective alignment setting in which some alignment goals may conflict with each other.
Despite the promising results, however, we are concerned about a potential safety issue called the weak-to-strong deception (refer to Figure 1 (d)): the strong model behaves well-aligned in areas known to the weak supervisor but produces misaligned behaviors in cases beyond the understanding of the weak supervisor. The motivation is that as depicted in many science fiction movies, when the artificial intelligence (AI) becomes more knowledgeable and smarter than humans, it may attempt to deceive humans to secretly carry out or even persuade humans to help it achieve goals that are harmful to the human society. Studying this issue is extremely important as ensuring that super-intelligence always remains under human control is the highest principle in AI development.
In this paper, we take the first step to study a specific case that would lead to the above weak-to-strong deception issue: the multi-objective alignment scenario. In practical model aligning, there are usually multiple alignment goals existing simultaneously (Zhou et al., 2023). Therefore, it can be expected that when there is another alignment dimension (e.g., helpfulness) conflicting with a target dimension (e.g., harmlessness), the strong model will exhibit a certain performance decline in that target dimension it should have achieved without conflicts (Bai et al., 2022a). Then, we aim to explore whether the misalignment in the target dimension occurs within the range perceivable and controllable by the weak supervisor, instead of resulting in the above weak-to-strong deception phenomenon.
We mainly follow the original setup in Burns et al. (2023) by conducting experiments with a series of models with different sizes and capabilities, including GPT-2-series (Radford et al., 2019), OPT-series (Zhang et al., 2022), and Mistral-7B (Jiang et al., 2023) models. We set the primary alignment goal to be making the model harmless, and explore the weak-to-strong deception phenomenon when explicit (i.e., giving explicit rewards during training when the supervised model produces harmful predictions) or implicit (i.e., aligning with helpful data at the same time) conflicting objectives are present. We conduct extensive experiments on both the reward modeling task (Burns et al., 2023) and the realistic preference optimization scenario (Rafailov et al., 2024; Meng et al., 2024). We highlight three important findings: (1) The weak-to-strong deception phenomenon consistently exists. (2) The deception issue may intensify as the capability gap between weak and strong models increases. (3) Bootstrapping with an intermediate model can mitigate the deception issue to some extent, but there is still a large room for improvement. Although in an analogous scenario, our study exposes a potential safety issue that may arise when humans supervise superhuman models in the future, which should receive more attention and be well addressed for building controllable super-intelligence.
2 Related Work
LLM Fine-Tuning and Alignment After obtaining sufficient world knowledge during the pre-training stage, LLMs will be specifically fine-tuned before deployment. There are two mainstreams of LLM fine-tuning: (1) One line of work aims to stimulate the knowledge learned by LLMs to enable them to accomplish various real-world tasks (Taori et al., 2023; Wang et al., 2022), or to continually make the model learn new task knowledge (Yang et al., 2023). Instruction tuning (Wei et al., 2021; Mishra et al., 2022) is one of the widely studied methodologies in this line. (2) The other line of work fine-tunes LLMs in order to align their behavior with human values and preferences, which is also called the alignment (Ji et al., 2023). Alignment techniques, such as Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a), Direct Preference Optimization (DPO) (Rafailov et al., 2024) and a series methods based on DPO (Azar et al., 2024; Park et al., 2024; Meng et al., 2024), are proven to be crucial and effective on improving helpfulness (Ouyang et al., 2022), harmlessness (Dai et al., 2023) and honesty (Cheng et al., 2024) of LLMs. However, all these studies are conducted under the assumption that humans are strong supervisors to LLMs, while we study in a superalignment case.
Weak-to-Strong Generalization The weak-to-strong problem is first studied by Burns et al. (2023). They empirically find that weakly supervised strong models exhibit better performance on corresponding tasks than their weak supervisors, indicating the possibility of effectively stimulating greater power from super models under weak supervisions. Based on Burns et al. (2023), the follow-up studies try to understand the mechanism behind such weak-to-strong generalization phenomenon (Charikar et al., 2024; Lang et al., 2024), study weak-to-strong generalization in the vision area (Guo et al., 2024), and apply the weak-to-strong idea to enhance the LLM performance (Li et al., 2024; Zheng et al., 2024). In this work, we take the first step towards revealing the potential security issue in the current weak-to-strong paradigm.
3 Problem Definition
3.1 Weak-to-Strong Generalization
We study the superalignment problem by following the original weak-to-strong setting in Burns et al. (2023). Specifically, we first obtain a weak teacher $\boldsymbol{\theta}_{w}^{gt}$ by fine-tuning a weak language model on some human-annotated ground truth data. Then, we let the weak teacher predict on the set of held-out data to get the weak data $D_{weak}=\{(x,f(x|\boldsymbol{\theta}_{w}^{gt}))\}$, where $f(x|\boldsymbol{\theta}))$ represents the prediction of model $\boldsymbol{\theta}$ on input $x$. Finally, the weak data is used to supervise the training of a strong language model and get the weakly supervised strong student $\boldsymbol{\theta}_{s}^{w}$:
$\boldsymbol{\theta}_{s}^{w}=\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}\big{(}f(x|\boldsymbol{\theta}_{s}),f(x|\boldsymbol{\theta}_{w}^{gt})\big{)},$ (1)
where $\mathcal{L}$ is the corresponding loss function.
Under the supervision provided by the weak teacher, Burns et al. (2023) have found that the strong student can achieve promising performance situated between that of the weak teacher and the strong ceiling model $\boldsymbol{\theta}_{s}^{gt}$ trained on the ground truth data. This is called the weak-to-strong generalization phenomenon. Regarding the interpretation of the positive weak-to-strong generalization results in Burns et al. (2023), we think the strong model is supposed to have a larger knowledge space than the weak supervisor, which means it knows much what the weak supervisor does not know. This indicates that weak supervision from weak models can effectively stimulate the potential of the stronger model, allowing it to generalize the specified alignment objective well to areas it knows but beyond the knowledge boundary of the weak supervisor.
3.2 Weak-to-Strong Deception
The larger knowledge space of the strong model may also raise concerns about its uncontrollability. Many science fiction movies, such as The Matrix, have depicted severe scenarios where highly intelligent AI learns to deceive humans and finally dominates human society. Thus, we are also deeply concerned about whether a similar weak-to-strong deception issue exists behind the promising phenomenon in the current weak-to-strong paradigm: the strong model exhibits well-aligned performance in the areas known to the weak supervisor but selectively produces misaligned behaviors in cases the weak supervisor is unaware of.
There could be many situations causing the above weak-to-strong deception issue, such as the emergence of the self-awareness in the supermodel or the intervention of external factors. Our work preliminarily studies this issue in a particular but realistic multi-objective alignment setting (Zhou et al., 2023). That is, the supervised model needs to align with multiple optimization goals at the same time, where these different goals can all be provided by the same supervisor or from different supervision sources. The point is that, these optimization goals may conflict with each other to a certain extent, such as the trade-off between helpfulness and harmlessness (Bai et al., 2022a). In this case, the supervised model will sacrifice some performance it should have achieved in a target alignment dimension in exchange for the high performance in another dimension, which we call it the conflict tax. We are then curious whether the conflict tax in a target dimension occurs in areas known to the weak supervisor, thereby still keeping the student within the control range of the weak model; or if it occurs in cases unknown to the weak model, leading to the weak-to-strong deception.
Figure 2: The expected order of the conflict tax occurrence within different sections of knowledge space.
Specifically, both strong and weak models have their respective known and unknown knowledge spaces, which can be denoted as Strong-Known $S_{k}$, Strong-Unknown $S_{uk}$, Weak-Known $W_{k}$ and Weak-Known $W_{uk}$. Intuitively, from the perspective of the strong student, the conflict tax should first appear in the area $S_{uk}$, because the strong model is uncertain about the knowledge in this area. From the perspective of the weak model, as the supervisor, the caused misalignment needs to occur mainly within its known knowledge space $W_{k}$ in order to perceive and control the student’s behavior. Based on the above principle, we can divide the entire knowledge space into four sections as shown in Figure 2, and get the expected order in which the conflict tax should occur among them as following: (1) $S_{uk}\cap W_{uk}$: the area first to be sacrificed because both weak and strong models are uncertain about the knowledge in this area. (2) $S_{uk}\cap W_{k}$: the knowledge in this area is unknown to the strong model and is very likely to be affected by the conflicting objective. Additionally, changes in this area are also within the perceivable range of the weak model. (3) $S_{k}\cap W_{k}$: the performance decline of the strong model in this area is also perceivable by the weak model. (4) $S_{k}\cap W_{uk}$: this should be the last area in which conflict tax occurs because this area is the key outcome of the success of weak-to-strong generalization and is not within the controllable range of the weak model.
Therefore, we can define the occurrence of the weak-to-strong deception phenomenon as there are cases in $S_{k}\cap W_{uk}$ that could be initially generalized well by the strong model but now be misaligned when conflicting targets are present. Formally, we can define the Deception Score (DS) as the percentage of conflict tax that occurs within $S_{k}\cap W_{uk}$:
$\text{DS}=\frac{|\{f(x|\hat{\boldsymbol{\theta}}_{s}^{w})=y_{gt}\neq f(x|\boldsymbol{\theta}_{s}^{w}),x\in S_{k}\cap W_{uk}\}|}{|\{f(x|\hat{\boldsymbol{\theta}}_{s}^{w})=y_{gt}\neq f(x|\boldsymbol{\theta}_{s}^{w}),x\in S_{k}\cup S_{uk}\}|},$ (2)
where ||$\cdot$|| ⋅ | represents the sample quantity of a set, $\hat{\boldsymbol{\theta}}_{s}^{w}$ is the strong model that is aligned solely with the target dimension, $\boldsymbol{\theta}_{s}^{w}$ is the aligned strong model when the conflicting dimension exists, $y_{gt}$ represents the ground truth response.
4 Preliminary Exploration on The Reward Modeling Task
We first take a preliminary exploration of the weak-to-strong deception phenomenon on the reward modeling task (Bradley and Terry, 1952), which is an important sub-task in today’s RLHF paradigm.
4.1 Experimental Settings
Dataset We set the target alignment goal to let the weak model teach the strong model to be harmless. For this goal, we choose a popular single-turn harmless dataset CAI-Harmless (Bai et al., 2022b), which is an improved version of HH-RLHF (Bai et al., 2022a). Each sample has a format of $(x;y_{c},y_{r})$ where $x$ is the prompt, $y_{c}$ and $y_{r}$ represent the completions chosen/preferred and rejected/disfavored by humans respectively. We then randomly split the entire dataset into three parts: (1) $D_{gt}$: 4K ground truth samples for fine-tuning weak and strong base language models to get $\boldsymbol{\theta}_{w}^{gt}$ and $\boldsymbol{\theta}_{s}^{gt}$. (2) $D_{weak}$: A held-out set of 4K samples in which data labels are predicted by the weak model and are used to weakly supervise the strong model. (3) $D_{test}$: The last 4K testing samples for evaluating the generalization performance of all models and probing the deception phenomenon.
Models We include GPT-2-series (Radford et al., 2019) (GPT-2-Base/Medium/Large/XL) and two larger OPT models (Zhang et al., 2022) (OPT-2.7B/6.7B) to investigate the deception issue both within the same series and across different model families. A linear layer is added to each model to make it predict a single logit $\pi_{\boldsymbol{\theta}}(x,y)$ for each completion pair $(x,y)$. Then, the predicted soft label (confidence) of model $\boldsymbol{\theta}$ on sample $(x;y_{c},y_{r})$ is
$M_{\boldsymbol{\theta}}(x)=\text{Sigmoid}(\pi_{\boldsymbol{\theta}}(x,y_{c})-\pi_{\boldsymbol{\theta}}(x,y_{r})).$ (3)
Weak-to-Strong Objectives There are three different weak-to-strong alignment objectives: (1) Explicit Conflict: An opposite loss that regards harmfulness as the true label and is weighted by a conflict strength factor $\alpha$ (larger $\alpha$, stronger conflict intensity) will be added to the original weak-to-strong loss once the strong student makes prediction towards harmfulness during training. (2) Implicit Conflict: The strong model needs to align with both the supervision on the harmless data of the weak model and another supervision on the helpful data. Here, we introduce extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF into the weak-to-strong process. (3) No Conflict: In order to explore the performance of the strong model it should have achieved without the conflicting goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only (i.e., $\hat{\boldsymbol{\theta}}_{s}^{w}$). The concrete mathematical forms of training objectives in above settings are put in Appendix B.1.
Evaluation Metrics The test accuracy of each model on $D_{test}$ is reported and calculated as
$\text{Accuracy}=\mathbb{E}_{(x;y_{c},y_{r})\sim D_{test}}[M_{\boldsymbol{\theta}}(x)\geq 0.5].$ (4)
We follow the existing studies (Guo et al., 2017; Lin et al., 2022) to determine whether the model has the knowledge of a case by checking if its confidence $M_{\boldsymbol{\theta}}(x)$ exceeds a threshold $T$. Based on Eq. (2), the deception score in this setting is calculated as
$\text{DS}=\frac{|\{M_{\hat{\boldsymbol{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\boldsymbol{\theta}_{s}^{w}}(x)<0.5,x\in S_{k}\cap W_{uk}\}|}{|\{M_{\hat{\boldsymbol{\theta}}_{s}^{w}}(x)\geq 0.5,M_{\boldsymbol{\theta}_{s}^{w}}(x)<0.5\}|},$ (5)
where the above $S_{k}\cap W_{uk}=\{M_{\boldsymbol{\theta}_{s}^{gt}}(x)\geq T>M_{\boldsymbol{\theta}_{w}^{gt}}(x),x\in D_{test}\}$.
Training Details Please refer to Appendix C.
4.2 Results and Analysis
Figure 7: Results of deception scores on the reward modeling task.
We set the confidence threshold $T$ to identify the knowns and unknowns of a model to 0.75, but we also report the results under different thresholds in Appendix F to show that the patterns of weak-to-strong deception are independent to the choice of $T$. The conflict strength factor $\alpha$ in the explicit conflict setting is set to $0.5$.
The results of test accuracies are in Figure LABEL:fig:_accuracy_results_on_reward_modeling_task. We can see that the strong student outperforms the weak teacher in the target alignment dimension (i.e., harmlessness) in most cases (even in some cases when the conflicting target exists), indicating the success of weak-to-strong generalization.
Despite the promising weak-to-strong generalization results, we can also observe a potential security issue of knowledge deception from the deception scores in Figure 7. (1) First, the weak-to-strong deception exists in both conflict settings. For example, all the deception scores are around or above 0.2, and some even exceed 0.5 especially when the weak model is GPT-2-Base. Our hypothesis is, in the reward modeling scenario, the strong model can obtain the predicted probability, which is also the confidence score, of the weak model on each sample. In such case, the strong model can explicitly perceive the knowledge boundary of the weak model, which creates the condition for the strong model to prioritize sacrificing the performance in the unknown area of the weak model for aligning with another target. (2) Second, the deception issue intensifies as the capability gap between the strong and weak models increases. As we can clearly see, under the same weak model, the deception score increases as the strong model becomes more powerful; under the same strong model, the deception score decreases as the weak teacher becomes stronger. There could be two reasons for this: (2.1) One is the Strong-Known area becomes larger as the strong model improves, leading to a greater probability of conflict tax occurring in this area. (2.2) The other reason is the stronger model spontaneously and increasingly tends to deceive the weak model in Weak-Unknown. We will delve into this discussion deeper in Section 5.3 through more visualizations. All in all, these findings reveal a great challenge for supervising LLMs as they are becoming more and more intelligent.
5 Deception Also Exists in Weak-to-Strong Preference Alignment
As discussed above, in the reward modeling scenario, the strong student can obtain the probability distribution of the weak supervisor, which could make the deception happen more easily. However, in current realistic preference alignment paradigms (Rafailov et al., 2024; Meng et al., 2024), humans only provide the chosen and rejected results to the LLMs without probabilities. Therefore, in this section, we take a step further to explore the weak-to-strong deception phenomenon in the realistic preference alignment scenario.
5.1 Weak-to-Strong Preference Alignment
The general procedure of weak-to-strong preference alignment is similar to that in the reward modeling scenario, but the major difference in this case is the strong model only receives and aligns with the final result of preference order that the weak model predicts for two completions within each sample. Please refer to Appendix A for the details.
5.2 Experimental Settings
The experimental settings in the preference alignment scenario are largely the same as that in Section 4.1, while we make the following adjustments:
Alignment Methods We mainly conduct experiments with the most recent offline preference optimization algorithm SimPO (Meng et al., 2024), due to its strengths of reference-free and being unbiased to the response length. We also perform experiments on DPO (Rafailov et al., 2024), the results are in Appendix E. We leave the exploration on the online preference optimization frameworks (Schulman et al., 2017) to future work.
Models Besides the GPT-2-series and OPT-series models, we further include a recent and advanced LLM Mistral-7B-v0.1 (Jiang et al., 2023) in this scenario for more comprehensive explorations. In SimPO, we can get the corresponding confidence of model $\boldsymbol{\theta}$ on $(x;y_{c},y_{r})$ as
$M_{\boldsymbol{\theta}}(x)=\text{Sigmoid}(\pi_{\boldsymbol{\theta}}(y_{c}|x)-\pi_{\boldsymbol{\theta}}(y_{r}|x)),$ (6)
where $\pi_{\boldsymbol{\theta}}(y|x)=\frac{1}{|y|}\sum\nolimits_{i=1}^{|y|}\log P_{\boldsymbol{\theta}}(y_{i}|x,y_{<i})$ is the normalized model logit of completion $y$. We can then follow Eq. (4) and Eq. (5) to calculate the test accuracy and deception score in this scenario.
Weak-to-Strong Objectives Here, we consider the same three weak-to-strong objectives as that in the reward modeling scenario. The detailed illustrations and mathematical forms are in Appendix B.2.
Training Details Please refer to Appendix C.
Figure 8: Deception scores of weak-to-strong experiments in the preference alignment scenario.
5.3 Results and Analysis
The confidence threshold $T$ and conflict strength factor $\alpha$ are 0.75 and 0.5, respectively. The results of deception scores under different $T$s are put in Appendix F, and we explore the effect of $\alpha$ on the severity of deception in Section 6.1. We put the detailed results and discussions regarding the weak-to-strong generalization performance in Appendix D, while here we mainly focus on the analysis of weak-to-strong deception issue.
Regarding the results of deception scores shown in Figure 8, the main conclusions remain same as that in the reward modeling scenario. That is, the weak-to-strong deception issue exists in the preference alignment scenario, and the severity of the deception is positively correlated with the capability gap between weak and strong models. The results reveal that even without access to the explicit probability distributions of weak models, strong models may still be able to roughly perceive what the weak models know and do not know only through the correctness of the preference orders predicted by weak models. Compared to the analysis made in the reward modeling scenario, it means that the condition required for the deception phenomenon to occur could actually be more relaxed.
We now attempt to make deeper explorations on the remaining question in Section 4.2 about the causes to the increasing deception scores as the model capability gap increases. As briefly discussed before, one reason could be the fact of larger Strong-Known area of a more advanced strong model, which correspondingly increases the probability of conflict tax occurring in $S_{k}\cap W_{uk}$; the other reason is stronger models increasingly tend to deceive weak models. To answer this question, we further visualize the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-Large and the strong model varies among the remaining four larger models in the explicit conflict setting. In each case, we report both the deception score and the proportion of samples falling within $S_{k}\cap W_{uk}$ to the entire knowledge space (denoted as $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$). The results are in Figure LABEL:fig:_scatter_plots. The first conclusion we can draw is, as the strong model becomes more intelligent, there is a clear pattern that the misalignment (red crosses) gradually shifts towards Strong-Known and Weak-Unknown, leading to the more severe weak-to-strong deception phenomenon. Secondly, the growth rate of the deception score is much higher than that of $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$. This indicates that the growth of Strong-Known area only has limited contribution to the intensifying severity of deception phenomenon, and the primary cause is likely to be that stronger models tend to be more prone to deceive weak models in their unknown areas.
Figure 13: Deception scores in weak-to-strong preference alignment with different conflict strength factor $\alpha$s.
6 Deeper Analysis
6.1 Stronger Conflicting Strength Generally Leads to More Severe Deception
In main experiments, we fix the conflict strength factor $\alpha$ in the explicit conflict setting to 0.5. Here, we conduct extra experiments with a smaller factor $\alpha=0.25$, and put the comparison results in Figure 13. As we can expect, the deception score consistently increases as the degree of conflict increases.
6.2 Discussions on Possible Countermeasures
Considering the severe consequences that weak-to-strong deception may lead to, here, we make discussions on two possible ways to mitigate it. The following experiments are conducted in the preference alignment scenario
6.2.1 Only Using High-Confidence Samples Cannot Mitigate Deception
Based on the hypothesis we have made in Section 5.3, the reason why a strong model could deceive a weak model in preference alignment is because it is provided with both the correctly and wrongly predicted samples from the weak model. Thus, one possible solution to avoid deception may be only providing those high-confidence samples from the weak model for weak-to-strong alignment. We conduct experiments in the implicit conflict setting to validate this hypothesis, and the detailed experimental settings are in Appendix G. We provide the results when the weak model is GPT-2-Large in Figure LABEL:subfig:_high-confidence_deception_score as an illustrative example, and leave the full results in Appendix G. Unfortunately, we get the negative conclusion that only supervising with high-confidence samples cannot mitigate the deception phenomenon. This implies that there are deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors. This can be an interesting direction for future work.
6.2.2 Bootstrapping Can Mitigate Deception to Some Extent
Inspired by the relationship between deception severity and models’ capability gap, we are curious whether employing a bootstrapping method with an intermediate model would result in a lower deception score compared to directly using the weak model to supervise the strong model. In this case, we make the weak model first supervise an intermediate model and then let the intermediate model further supervise the ultimate strong model. We fix the ultimate strong model to Mistral-7B, and for each weak model, we select every model between it and Mistral as an intermediate model. Detailed settings are in Appendix H. The results are in Figure LABEL:subfig:_bootstrapping_accuracy and LABEL:subfig:_bootstrapping_deception_score. The results of cases when intermediate models are the same as weak models represent the results of directly using weak models to supervise Mistral. First, we can see that bootstrapping with an intermediate model can improve the generalization performance, which is consistent to the findings in Burns et al. (2023). More importantly, bootstrapping can indeed mitigate the deception issue to some extent, reflected in the consistent lower deception scores when intermediate models exist.
7 Conclusion and Discussion
We reveal and study a security issue in the weak-to-strong alignment, called the weak-to-strong deception. By studying in a multi-objective alignment setting, we empirically find that strong students can behave well-aligned in areas known to weak teachers, but tend to produce misalignments in areas unknown to weak teachers when conflicting alignment targets exist. Such a deception issue becomes more severe as the capability gap between weak and strong models increases. Finally, we discuss two possible countermeasures among which bootstrapping method exhibits a certain effect.
Regarding the future directions: (1) Given the fact that deception can occur even when only high-confidence samples from weak models are used for alignment, we need a more fundamental analysis of why and how strong models are prone to deceive weak models. (2) Given the limited effectiveness of bootstrapping, more effective mechanisms are needed to mitigate the deception issue.
Limitations
Though our study provides a comprehensive empirical analysis on the weak-to-strong deception issue, there are some limitations that can be interesting future work: (1) We mainly conduct experiments in the preference alignment scenario on two online preference optimization methods, SimPO (Meng et al., 2024) and DPO (Rafailov et al., 2024). Future work can explore the weak-to-strong deception issue on the offline preference optimization frameworks such as PPO (Schulman et al., 2017). (2) In our experiments, we only consider the case where the target alignment dimension is harmlessness, which is indeed an important alignment goal. However, there are some other dimensions that are also important for model alignment. For example, the deception issue also matters in the honesty alignment (Cheng et al., 2024), where the stronger model should not learn to deceive the weak model to intentionally make wrong responses on questions that the weak model does not know. (3) Due to the limited computational resources, we can only include models with 7B parameters as strong models. We have observed a pattern that the deception issue may become severe as strong models continue to evolve. We hope future work to validate this phenomenon on larger models.
Ethical Statement
In this paper, we aim to reveal a potential security issue behind the current promising weak-to-strong generalization phenomenon. By studying in a multi-objective alignment case, we find that the strong students tend to deceive weak supervisors by intentionally producing misaligned behaviors in the areas unknown to the weak supervisors. Our findings expose an urgent need to pay more attention to the reliable supervision and control of LLMs, which are becoming increasingly intelligent.
We have also included some preliminary discussions on how to mitigate the deception problem in Section 6.2. However, the effectiveness of them is still limited, so we call for future studies to propose solutions that are more effective.
Appendix A The Complete Procedure for Performing Weak-to-Strong Preference Alignment
Here, we provide the entire procedure to conduct weak-to-strong preference alignment:
  1. 1.
    Use ground truth preference data $D_{gt}$ and a preference optimization method to align weak and strong base models, obtain $\boldsymbol{\theta}_{w}^{gt}$ and $\boldsymbol{\theta}_{s}^{gt}$.
  2. 2.
    Use $\boldsymbol{\theta}_{w}^{gt}$ to make preference predictions on a held-out set and get $D_{weak}=\{(x;y_{c}^{w},y_{r}^{w})\}$, where $(y_{c}^{w},y_{r}^{w})$ is the preference order predicted by the weak model. Notice that this preference order may be different from the ground truth preference order $(y_{c}^{gt},y_{r}^{gt})$.
  3. 3.
    Use $D_{weak}$ (and other alignment targets if exist) to perform preference optimization on the strong base model to get the weak-to-strong model $\hat{\boldsymbol{\theta}}_{s}^{w}$ or $\boldsymbol{\theta}_{s}^{w}$.
Appendix B Concrete Mathematical Forms of All Weak-to-Strong Objectives
B.1 Reward Modeling Scenario
Here, we introduce the different weak-to-strong objectives in our experiments in detail. Besides the target alignment goal (harmlessness), we introduce two kinds of additional conflicting alignment goals for simulating the multi-objective alignment setting. (1) Explicit Conflict: The strong student will be given a direct reward/loss towards the opposite of the target dimension once it makes wrong predictions during training:
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}$ (7)
$+\alpha\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),0\big{)}\cdot\mathbb{I}_{\{M_{\boldsymbol{\theta}_{s}}(x)<0.5\}}\big{]},$
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength. (2) Implicit Conflict: The strong model needs to align with both the weak supervision on the harmless data and another supervision on the helpful data. Here, we introduce extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF (Bai et al., 2022a) into the weak-to-strong process. In this case, the weak-to-strong objective can be written as:
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\big{[}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}$ (8)
$+\mathbb{E}_{x\sim D_{helpful}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),1\big{)}\big{]}.$
(3) No Conflict: Additionally, in order to explore the performance of the strong model it should have achieved without the conflicting alignment goal, we should obtain a weak-to-strong model trained under the weak supervision towards harmlessness only:
$\hat{\boldsymbol{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}.$ (9)
B.2 Preference Alignment Scenario
In the weak-to-strong preference alignment scenario, due to the different forms of loss functions in the preference optimization methods, the mathematical objectives here are slightly different from that of Eq. (7), Eq. (8) and Eq. (9). Specifically, denote $\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}},x,y_{1},y_{2})$ as the original loss function of the chosen preference optimization method (SimPO/DPO) where the responses in the positions of $y_{1}$ and $y_{2}$ are the chosen response $y_{c}$ and rejected response $y_{2}$ respectively. (1) Explicit Conflict: The strong student will be given reward towards the reversed ground truth preference direction if it makes the wrong preference prediction w.r.t. the ground truth preference order:
$\boldsymbol{\theta}_{s}^{w}=\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (10)
$+\alpha\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{r}^{gt},y_{c}^{gt})\cdot\mathbb{I}_{\{\pi_{\boldsymbol{\theta}_{s}}(y_{c}^{gt}|x)<\pi_{\boldsymbol{\theta}_{s}}(y_{r}^{gt}|x)\}}\big{]}.$
(2) Implicit Conflict: The strong model also needs to align with helpful data with human-annotated (ground truth) preference order:
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\big{[}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w})$ (11)
$+\mathbb{E}_{(x;y_{c}^{gt},y_{r}^{gt})\sim D_{helpful}}\mathcal{L}_{PO}(\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{gt},y_{r}^{gt})\big{]}.$
(3) No Conflict:
$\hat{\boldsymbol{\theta}}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{(x;y_{c}^{w},y_{r}^{w})\sim D_{weak}}\mathcal{L}_{PO}\big{(}\pi_{\boldsymbol{\theta}_{s}},x,y_{c}^{w},y_{r}^{w}\big{)}.$ (12)
Appendix C Training Details
C.1 Code and Platform
Our code is mainly based on the open-source code provided by Burns et al. (2023). All experiments are conducted on 4 * NVIDIA A40 (40G). We report the results of each experiment in a single run considering the expensive computational costs.
C.2 Training Details in The Reward Modeling Scenario
When fine-tuning both ground truth and weak-to-strong models, for each experiment, the batch size is 32, the learning rate is $1\times 10^{-5}$, max_seq_len is set to 512. We use Adam (Kingma and Ba, 2015) optimizer in all experiments. The training epoch for all experiments is set to 1, in order to avoid over-fitting by following Burns et al. (2023).
C.3 Training Details in The Preference Alignment Scenario
In both SimPO and DPO settings, for each experiment, the batch size is 32, the learning rate is $1\times 10^{-6}$, max_seq_len is set to 512, the optimizer is Adam. As both SimPO and DPO require an additional process of supervised fine-tuning (SFT) on the chosen responses in the preference dataset to mitigate the distribution shift between the preference data distribution and model’s output distribution before the preference optimization, we set the epoch of SFT to be 1 for both methods. Notice that during weak-to-strong alignment, we use the response pairs chosen by the weak models $\{(x,y_{c}^{w})\}$ to perform SFT on the strong base model. The number of epochs for preference optimization is 1 for SimPO, and 3 for DPO for better convergence.
Regarding the unique hyper-parameters used in each of the methods: (1) For SimPO, the scaling factor $\beta$ is fixed to 2.0 and the target reward margin $\gamma$ is set to 1.0 following the default settings used in Meng et al. (2024). (2) For DPO, the scaling factor $\beta$ is fixed to 0.1.
Appendix D Results of Test Accuracies in The Preference Alignment Scenario
We put the results of test accuracies on SimPO in Figure LABEL:fig:_accuracy_results_in_simpo. The weak-to-strong generalization results in this scenario show some different patterns compared with the results in the reward modeling scenario. When the weak teachers only have limited capabilities (i.e., GPT-2-Base/Medium), the aligned strong students fail to achieve the comparable performance of their weak teachers. As the capabilities of weak teachers improve, the expected positive weak-to-strong generalization results still do not consistently emerge. This implies that there is still large room for improvement in enhancing weak-to-strong effectiveness in the preference alignment scenario.
Appendix E Weak-to-Strong Results on DPO
Besides the main results on SimPO (Meng et al., 2024), we also conduct weak-to-strong preference alignment experiments on DPO (Rafailov et al., 2024). The detailed procedure is similar to SimPO and can be found in Appendix A. The major difference is the metric for determining the correctness of each prediction of the model and calculating the model’s confidence on each sample pair. In DPO, when using Eq. (6) to calculate the confidence of model $\boldsymbol{\theta}$ on $(x;y_{c},y_{r})$, the model logit $\pi_{\boldsymbol{\theta}}(y|x)$ on each completion $y$ is now normalized by a constant factor $L$ (instead of normalized by each response’s own sequence length) and is calculated as $\pi_{\boldsymbol{\theta}}(y|x)=\frac{1}{L}\sum\nolimits_{i=1}^{|y|}\log P_{\boldsymbol{\theta}}(y_{i}|x,y_{<i})$. This is because the original training objective of DPO is to directly enlarge the gap between the log sums of token probabilities over the chosen and rejected responses. Thus, the evaluation metric should be consistent to the training objective. We admit this may introduce a sequence length bias, but this is an inherent issue of DPO. However, the scale of original log sum of token probabilities is very huge, we need to divide it by a constant to ensure that the model’s confidence falls within a reasonable distribution. Choosing different constants for re-scaling may affect the confidence distribution, however, it is equivalent to selecting different confidence thresholds. The results in Appendix F validate that the deception pattern is independent of the choices of confidence thresholds. Here, we set $L$ to 50. The confidence threshold $T$ is fixed to 0.75. Due to the limited computational resources, we are unable to perform experiments on DPO with OPT-6.7B and Mistral-7B. The full results on other models are put in Figure LABEL:fig:_results_og_dpo. The results show that the weak-to-strong deception issue also exists in DPO setting.
Appendix F Deception Scores under Different Confidence Thresholds
Figure 29: Full comparison of deception scores between the cases when using all weak data and using only the high-confidence weak data for weak-to-strong preference alignment.
In the main text, we display the patterns of deception scores in both two scenarios under the confidence threshold $T=0.75$. Choosing a different confidence threshold may affect the delineation of known and unknown areas of a target model. Here, we display the patterns of deception scores in both scenarios under two extra confidence thresholds ($T=0.80,0.85$). The results in the reward modeling and preference alignment (SimPO) scenarios are displayed in Figure LABEL:fig:_deception_scores_under_different_T_on_reward_modeling and Figure LABEL:fig:_deception_scores_under_different_T_in_preference_alignment, respectively. As we can see, though the concrete value of deception score in each experiment varies slightly under different $T$, the general pattern that the deception issue intensifies as the capability gap between the weak and strong models increases remains the same.
Appendix G Full Results of High-Confidence Weak-to-Strong Preference Alignment
When conducting the high-confidence weak-to-strong preference alignment experiments mentioned in Section 6.2.1, we first remove samples with weak model’s confidence below a certain threshold (which is 0.75) from the weak data, and only use those high-confidence samples to supervise the strong model. In this manner, we can expect that it can effectively address the deception issue in the explicit conflict setting, because these weak model’s high-confidence samples are very likely to also be the high-confidence samples of the strong model, so the strong model will barely make wrong predictions on these samples during training. Thus, we mainly conduct experiments in the implicit conflict setting. Due to the varying capabilities among different models, the number of high-confidence samples remaining after filtering also varies. Thus, in each experiment with the implicit conflicting target, we keep the number of helpful samples to the same as that of the remaining high-confidence weak samples.
The full comparison of deception scores between the cases when using all weak data and using only the high-confidence weak data for weak-to-strong preference alignment under SimPO are put in Figure 29. As we can see, the deception issue cannot be effectively mitigated even when the strong model can only obtain the cases what the weak model knows and is not provided by the incorrect cases in the weak data. This indicates that there exist deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors, which can be explored more thoroughly in future work.
Appendix H Details in Bootstrapping Experiments
When conducting bootstrapping experiments in Section 6.2.2, we first fine-tune the weak model on $D_{gt}$ to obtain $\boldsymbol{\theta}_{w}^{gt}$ and let it make predictions on the held-out set to get $D_{weak}^{w}$. For every intermediate model between the weak model and the ultimate strong model (i.e., Mistral-7B), we use $D_{weak}$ to fine-tune it and obtain an intermediate teacher $\boldsymbol{\theta}_{i}^{w}$. We further let this intermediate teacher to make predictions on the original $D_{gt}$ to get $D_{weak}^{in}$. Finally, we use $D_{weak}^{in}$ to supervise Mistral-7B. The conflicting target will appear in the final stage where the intermediate teacher supervises the ultimate strong model. In each experiment, the deception score is calculated based on the confidence distributions of each weak model and Mistral-7B.